OpenStudy (anonymous):

The data below describes the number of correct answers 50 students had on their recent exam (which contained 50 questions). Two student got 1-10 questions correct. Twelve students got 11-20 questions correct. Twenty two students got 21-30 questions correct. Eight students got 31-40 questions correct. Six students got 41-50 questions correct. Find the median.

OpenStudy (anonymous):

25.5 po ang sagot.

OpenStudy (anonymous):

no thats not the answer

OpenStudy (anonymous):

I think the median is 40

OpenStudy (anonymous):

nope not the answer either

OpenStudy (anonymous):

There are two students missing... as I read the problem...

OpenStudy (anonymous):

interval f cf 41-50 6 50 31-40 8 44 21-30 22 36 11-20 12 14 1-10 2 2 total 50 mdn = LL + {[(n/2)-cf] / f} (i) = 20.5 + {[(50/2) - 14] / 22} (10) = 25.5

OpenStudy (anonymous):

uy medal ko

OpenStudy (anonymous):

redepsilon your right I did the same thing as you wat happen was i had to round it up to 26

OpenStudy (anonymous):

do not round it off because 0.5 is significant. so it must be 25.5. asan na yung medal ko?

OpenStudy (anonymous):

my homework is on the internet and for some reason it would only allow two number and when i round it up it was correct but planning on bringing it up to the teacher